A stone was thrown vertically upward and it rised 10m above its initial location before it 

started to fall. What was its initial velocity?


1
Expert's answer
2021-06-14T13:59:07-0400

According to the energy conservation law, the initial kinetic energy (at the start point) of the stone is equal to its potential energy at the highest point of its trajectory:


"\dfrac{mv^2}{2} = mgh"

where "m" is the mass of the stone, "v" is its initial velocity, "h = 10m" is the highest heigth, "g = 9.8N\/kg" is the gravitational acceleration. Expressing "v", obtain:


"v = \sqrt{2gh}\\
v = \sqrt{2\cdot 9.8\cdot 10} = 14m/s"

Answer. 14 m/s.